OpenStudy (anonymous):

Stuck again. Find x. 2 1 11 --- + --- = --- x 3 3x

OpenStudy (anonymous):

I hope that's clear enough. (2)/x + (1)/3 = (11)/3x

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

2 1 11 --- + --- = --- multiply both sides by x to get x 3 3x 2+ x/3 = 11/3 x/3=11/3 -2 =5/3 x=3(5/3) =5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

I don't understand the 2nd step, mark o.

OpenStudy (anonymous):

x/3 = 11/3 - 2 = 5/3

OpenStudy (anonymous):

2+ x/3 = 11/3 x/3=11/3 -2 =11/3 -6/3=5/3 x=3(5/3) =5

OpenStudy (anonymous):

is it better now OMG

OpenStudy (anonymous):

Yes, it makes more sense! And the answer should be 5. Thanks for showing the work!

OpenStudy (anonymous):

wc n thnks